It is a function of x and y. and x reoso, y = rsin. give Vf. VL. ||||2 in both rectangular and polar coordinates,
Answer #1

5:07 AM Vo 0.OKB/s 90 KEY TAKEAWAYS Key Points To convert from polar to rectangular (Cartesian) co- ordinates use the followiVo 5:17 AM 0 0.0KB/s 90 7.3 Polar Coordinates WN a Using right-triangle trigonometry, the following equations are true for thVo 5:32 AM 2.0KB/s 87 Х How to obtain the gradi... https://math.stackexchange.com a The gradient operator in 2-dimensional CaVo 5:33 AM 80.8KB/s 87 а де Ә — + дх д ду От 0 ах дү де о дах дө дө Ә + ду де ду дө Observing that r = х2 +y2 and 0 = arctan(Vo 5:33 AM 0.0KB/s 87 This certainly gives the right answer, but there is a quicker way. Consider a function f(r) in polar coVo 5:33 AM 0.0KB/s 87 Observe that d f = dr. Vf and dr = d rêr +rd deo. Suppose, then, that Vf = aềr + ßeo where a and ß areg n and yersino sol ܢ 2 yersino 7 Civen fis hinction of (se and narcoso To hind ® gf, vt, llyf? Given nex-coso, yersing fin,yogn ulas (2) where f is hinction of me and y we have If (119) 2 tinig); + 2 finly) ? he or fa at at of a Î bi he for poloar c5:19 AM 0.OKB/s & Neul 89 Google Q what is the gradient of f(x y) ALL VIDEOS IMAGES NEWS SHOPPING The gradient of a function,5:40 AM 0.1KB/s Vo LTE 87 = div Vf = V.Vf a a a it j + k Əx ду дz af af af j+ k Əx ду Əz a af a af a + Əx дх Dy ( Dy az ²f aftо е Give thumps up conclusion for pdan Coordinates + Нео)* 2 24 ée || or | # 22-2. (ъ)*- ә27 0 ify eg 22-р Deore De ee h әү